OpenStudy (anonymous):

Find the centroid of triangle EFG with vertices at E(-5,5), F(-1,3), and G(-3,1) How do i find the answer to this problem?

OpenStudy (anonymous):

Use two medians and find their intersection point.

OpenStudy (anonymous):

midpoint M of EF = (-3, 4) slope of median GM = 0 y - 4 = 0(x- (-3)) y=4

OpenStudy (anonymous):

midpoint N of FG = (-2, 2) slope of median EN = 3/7 y - 2 = 3/7(x-(-2)) y = (3/7)x + 6/7+ 2 y = (3/7)x +20/7

OpenStudy (anonymous):

Now just solve for the intersection of these two medians.